A gas has a volume of 800 mL at 23°C and 300 torr. What would the volume of the gas be at 227°C and 600 torr of pressure?

1
Expert's answer
2022-03-13T15:20:18-0400

P1V1÷T1=P2V2÷T2

P1=300 torr

V1= 800ml

T1= 23+273=296K

P2=600 torr

V2=?

T2=227+273=500K

300×800÷296 = 600V2÷500

V2= 300×800×500÷(296×600)

=675.68ml